The investors of Smith Autos have put up $500,000 to construct a building and purchase all equipment required to wash cars. The investors expect to earn a minimum return of 10% on their investment. If the money to set up the business had been borrowed from a bank instead, the car wash owners would have paid a 10% interest rate. The car wash is open 50 weeks per year and washes 800 cars per week. Whether operative or not, the company must pay $1,000 per week return to investors and $1,000 per week as insurance. The variable costs for the 800 weekly washes includes $1,000 labour cost and $600 materials cost. There are many car washes of equal quality and service in the area and they charge $5 per car wash.a) Based on your knowledge of the different forms of market structure, do you think Smith Autos should stay in business? Show calculations. (2 Marks)b) Graphically represent the company's performance, showing profit or loss. (1 Mark)c) Another car wash company operates inside a gated community in the vicinity and offers its services at $6 per wash under a licensed agreement. Highlight how is it possible for this company to charge a higher price than the market and still stay in business. (1 Mark)After a few mouths, due to the forces of demand and supply, the market price is pushed down to $1.5 per car wash. What course of action will you suggest to Smith Autos? (1 Mark

1
Expert's answer
2021-01-26T08:49:43-0500

Q = 800 cars per week,

FC = 1,000 + 1,000 = 2,000 per week,

VC = 1,000 + 600 = $1,600,

AVC = VC/Q = 1,600/800 = $2.

ATC = (FC + VC)/Q = $4.5.

P = $5 per car wash.


a) This market structure is perfectly competitive, so as P > AVC and P > ATC (5 > 4.5), then Smith Autos is profitable and should stay in business.


b) Graphically the company's performance can be represented as P = MC above the ATC curve, so its profit is a rectangle with area (P - ATC)×Q.


c) If another car wash company operates inside a gated community in the vicinity and offers its services at $6 per wash under a licensed agreement, then it can operate as a monopoly, so it is possible for this company to charge a higher price than the market and still stay in business.

If after a few months, due to the forces of demand and supply, the market price is pushed down to $1.5 per car wash, then Smith Autos should shut down, because P < AVC.